# The Compilation and Validation of the Spectroscopic Redshift Catalogs for the DESI-COSMOS and DESI-XMMLSS Fields

### v2

@@ Appendix C Data Model
DESINAME |S22 Unique DESI target identifier
@@
PMRA mas / yr >f4>f4 Proper motion in the +RA direction (already including cos(dec))
PMDEC mas / yr >f4>f4 Proper motion in the +Dec direction
@@
LINENAME_FLUX $1\times 10^{-17}\mathrm{erg}/(\mathrm{s}\,\mathrm{cm}^{2})$ >f4>f4 Gaussian-integrated emission-line flux measurement from FastSpecFit of the best spectrum
LINENAME_FLUX_IVAR $1\times 10^{34}\mathrm{cm}^{4}\mathrm{s}^{2}/\mathrm{erg}^{2}$ >f4>f4 Inverse variance of integrated flux measurement from FastSpecFit of the best spectrum
@@
N_EPOCHS Number of different observational epochs for the target
LRG_MASK >f8>f8 Bright star mask used for LRG targets applied to the catalog
ELG_MASK >f8>f8 Bright star mask used for ELG targets applied to the catalog
@@
VRAD km / s >f8>f8 Radial velocity
VRAD_ERR km / s >f8>f8 Radial velocity error
@@
VSINI km / s >f8>f8 Stellar rotation velocity

### v3

@@ Appendix C Data Model
DESINAME |S22 Human readable identifier of a sky location DESI JXXX.XXXX[+/]$Y$.YYYYJXXX.XXXX[+/-]$Y$.YYYY, where X,Y=truncated decimal TARGET_RA, TARGET_DEC, precise to 0.36 arcsec. Multiple objects can map to a single DESINAME if very close on the sky.
@@
PMRA $\rm{mas}\,yr^{-1}$ >f4>f4 Proper motion in the +RA direction (already including cos(dec))
PMDEC $\rm{mas}\,yr^{-1}$ >f4>f4 Proper motion in the +Dec direction
@@
LINENAME_FLUX $10^{-17}\mathrm{erg}/(\mathrm{s}\,\mathrm{cm}^{2})$ >f4>f4 Gaussian-integrated emission-line flux measurement from FastSpecFit of the best spectrum
LINENAME_FLUX_IVAR $10^{34}\mathrm{cm}^{4}\mathrm{s}^{2}/\mathrm{erg}^{2}$ >f4>f4 Inverse variance of integrated flux measurement from FastSpecFit of the best spectrum
@@
LRG_MASK $uint8$ Bright star mask used for LRG targets applied to the catalog
ELG_MASK $uint8$ Bright star mask used for ELG targets applied to the catalog
@@
VRAD $\rm{km}\,s^{-1}$ >f8>f8 Radial velocity
VRAD_ERR $\rm{km}\,s^{-1}$ >f8>f8 Radial velocity error
@@
VSINI $\rm{km}\,s^{-1}$ >f8>f8 Stellar rotation velocity